OpenStudy (anonymous):

Find f(1), f(2), f(3), and f(4) for the function represented by each rule. f(n)= -2n

OpenStudy (anonymous):

f(1), plug in 1 for n in the function f(n): f(1) = -2(1) f(1) = -2 Do this for the rest of the values: f(2) = -2(2) = -4 f(3) = -2(3) = -6 f(4) = -2(4) = -8